OpenStudy (anonymous):

Please help me =*( A certain taxi charges $0.85 for the first ½ mile and $0.25 for every ½ mile after that. The total cost of a trip was $8.85. Which one is greater? A. The trip's distance in miles B. 16

OpenStudy (anonymous):

i got 32.5

OpenStudy (anonymous):

correction

OpenStudy (anonymous):



OpenStudy (anonymous):

with 1 dollar it covers=2 miles with 8 dollars it covers=2*8=16 miles distance covered\[=16+\frac{ 1 }{ 2 }=16\frac{ 1 }{ 2 }=16.5 ~miles\]

OpenStudy (anonymous):

so my method is wrong?

OpenStudy (anonymous):

i have to eat brb

OpenStudy (anonymous):

x is no. of half miles .85+.25(x-1)=8.85 .25x-.25=8.85-0.85=8 0.25x=8+.25=8.25 x=\[\frac{ 8.25 }{ 0.25 }=\frac{ 825 }{ 25 }=33\] 33 half miles=16.5 miles

OpenStudy (anonymous):

hey so did i do it right?

OpenStudy (anonymous):

please note the correction.

OpenStudy (anonymous):

it is x-1 not x-0.5

OpenStudy (anonymous):

:)